Scores of an IQ test have a​ bell-shaped distribution with a mean of 100 and a standard deviation of 14. Use the empirical rule to determine the following. ​(

a) What percentage of people has an IQ score between 72 and 128​?

​(b) What percentage of people has an IQ score less than 58 or greater than 142​? ​

(c) What percentage of people has an IQ score greater than 128​?

- Type as an integer or decimal
Answer #1

Ams) Given data Mean ll =100 - standard deviction & =14 According to emperical role... within i standard deviation there is aatcr c) Percentage of the people greater than 128 P(x>128) = P(Z > 1287400) = P(Z >2) El-p(262) = 1-0.977 = 0.023 2.3% of peo